A car started out from Memphis toward Little Rock at the rate of 60 km/h. A second car left from the same point 2h later and drove along the same route at 75km/h. How long did it take the second car to overtake the first car?
Answer by Paul(988)

A car started out from Memphis toward Little Rock at the rate of 60 km/h. A second car left from the same point 2h later and drove along the same route at 75km/h. How long did it take the second car to overtake the first car?
Let the time be x
60(x+2)=75x
60x+120=75x
60x+120-60x=75x-60x
120=15x (divide both sides by 15)
x=8